Solve the following equation for v:
S = (-1/2)gt^2 + vt
subtract (-1/2)gt^2 from both sides
s - (-1/2)gt^2 = vt, or,
s + (gt^2/2) = vt
then divide by t
s/t + gt^2/2t = v, or,
s/t + gt/2 = v
therefore, v = s/t + gt/2

Answer by geochica(3)

the original equation:
S = (-1/2)gt^2 + vt
we want to isolate "v". so we need to cancel out everything else on the right side of the equation.
subtract (-1/2)gt^2 from both sides
s - (-1/2)gt^2 = vt, or,
s + (gt^2/2) = vt
then divide by t
s/t + gt^2/2t = v, or,
s/t + gt/2 = v
therefore, v = s/t + gt/2